If a discount loan has a face value of $4200 and yields proceeds of $3759, find the amount of discount interest charged.
Answer by Theo(13342)

with a discount loan, the interest is subtracted from the principal.
the borrower gets the net value of the original loan amount minus the interest.

in this problem, the face value of the loan is 4200.

the loan yields proceeds of 3759.

this, i presume, is what the borrower gets.

if correct, that means the interest charged is 4200 - 3759 = 441.

the interest was calculated off the 4200, i believe.
that would be 441 / 4200 = .105 = 10.5%

the effective interest would be 441 / 3759 = .1173184358 = 11.73% rounded to two decimal places.

i believe your solution is that the amount of discount interest charged is 441.

that's the difference between the face value of the loan and what the borrower gets.

the borrower gets 3759, but has to pay back 4200.

here's a reference on discount loans.

https://marketbusinessnews.com/financial-glossary/discount-loan/